### v13

@@ Abstract
In an exchange economy composed of risk-averse individuals facing shocks, these shocks may become uninsurable if they are private information. Efficient shock smoothing across the economy requires self-reporting of the shocks that each individual privately suffers. This efficient allocation of resources leads to growing inequality between individuals, and the problem remains largely unsolved. This paper proposes a solution. The proposed mechanism (i) promises within-period full insurance by postponing risks ; (ii) does not lead to social ranking among the same age groups; and (iii) is sustainable under some conditions.
Key words: Incentives, optimal resource allocation, recursive contract, equality of same age group.
@@ 1 Introduction
In an exchange economy composed of risk-averse individuals facing shocks, these shocks may become uninsurable if they are private information. This problem has been studied extensively since the work of Green [1987]. These shocks are uninsurable because efficient shock smoothing across the economy requires self reporting of shocks that each individual privately suffers. Proposed solutions are limited. Inducing truth-telling at the expense of equality leads to immiserization, where the average expected lifetime utility of individuals converge to negative infinity while driving inequality without bound[Green, 1987, Thomas and Worrall, 1990]. If discount factor is approximately one, meaning individuals value the future and care little about the present, their contracts approach to the first best [Thomas and Worrall, 1990, Carrasco et al., 2019]. Atkeson and Lucas [1992] consider a scenario where discount factors represent resource prices, and social planners trade claims to current and future resources at these prices. They show that full information is required for efficient allocation.
This paper extends an alternative mechanism proposed by Marcet and Marimon [1992], which distributes risk across periods for the same individual. This approach could reduce the inequality observed in previous studies that allocate risk among individuals within each period. The mechanism proposed by Marcet and Marimon [1992] was an efficient growth mechanism combining state-contingent investment and transfers. This paper uses it for transfers in an exchange economy facing income shocks.
The first key finding of this paper is that our mechanism can achieve both efficiency and avoidance of immiserization under certain conditions   specifically, when individuals’ degree of risk aversion is not excessively high relative to the disparity between their highest and lowest possible incomes. Furthermore, the mechanism preserves the equality of opportunity within the same age group. Both efficiency and equality of opportunity are achieved by shifting risks to future periods. However, if the degree of risk aversion is too high relative to the disparity in income levels, challenges may arise. In such cases, it may become necessary to temporarily adjust the optimal transfer rules. This disruption occurs when the promised lifetime utility cannot be defined, given the assumption that each individual’s utility function is bounded from above. Nevertheless, even when the sufficient condition is not met, the mechanism continues to maintain the optimal allocation under the given condition and avoids immiserization.
The second key finding pertains to the potential sustainability of this mechanism in a society with a continuum of individuals, where the population, consisting of different age group, remains constant over time. In such a society, each individual faces a probability of dying in the next period, while new individuals continuously enter. Since risks exist across periods, maintaining unchanged expected lifetime utility across periods incurs costs. To ensure sustainability, earlier entrants must bear these costs, resulting in average consumption increasing with age. However, this does not lead to the formation of social ranking within the same age group. Importantly, this mechanism can ensure higher expected lifetime utility compared to autarky for all generations, while sustaining a positive balance of payments over time. Thus, a society that achieves both efficiency and equality of opportunity is feasible.
The paper is structured as follows. Section 2 introduces the notation and tarminology and proposes the mechanism for the inter-period risk transfers. Section 3 shows efficiency of the proposed mechanism and examines the properties for individuals. Section 4 investigates the potential sustainability of the mechanism. Section 5 provides some numerical examples. Section 6 evaluates the mechanism’s efficiency again using Bellman equation. Section 7 summarizes the conclusions.
@@ 2 Model
Consider an infinitely existing risk-neutral government for a society with a continuum of individuals in a unit interval. Each individual is risk-averse and will be alive in the next period with probability $\alpha\in(0,1)$. In each period, newly born risk-averse individuals enter the society, holding the total population constant. Individuals differ only in their periods of birth and death. Each individual has a discount factor $r\in(0,1)$. Let $\beta=\alpha r$ be the multiplication of $r$ and $\alpha$. A similar stationary population model is used in Fujiwara-Greve and Okuno-Fujiwara [2009] for their Prisoner’s Dilemma game. Let $(\Omega,\mathcal{F},P)$ be a probability space, and let $\{e_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ be a sequence of integrable random variables on $(\Omega,\mathcal{F},P)$ denoting identically and independently distributed income shocks of each individual in a sequence of periods. These income shocks take values in $E:=\{e^{1},e^{2},\dots,e^{M}\}$, where $e^{i}\in\mathbb{R}_{+}$ with $e^{i}<e^{j}$ for $i<j\in\{1,\dots,M\}$, $M\geq 2$. The government considers a transfer mechanism $\mathcal{M}$ that repeats from period zero to infinity, using a sequence of transfers (controls) $\{\tau_{t}\}_{t=0}^{\infty}$, $\tau_{t}\in\mathbb{R}$ and a sequence of utility control variables $\{\lambda_{t}\}_{t=0}^{\infty}$, $\lambda_{t}\in\mathbb{R}_{++}$ which will be specified later.
2.1 Pareto optimal transfer mechanism with a utility control variable $\lambda_{0}$
@@
First, we derive an optimal transfer function which, under symmetric income information, ensures a level of utility corresponding to a utility control variable $\lambda_{0}>0$ Following Marcet and Marimon [1992], we will use this utility control variable $\lambda_{0}$ as a state variable under asymmetric income information. For this purpose, in this subsection only, we consider the case where individuals cannot hide their income from the government in each period.
@@
In period 00, the government is considering an efficient transfer mechanism that increases individuals’ lifetime utility as much as possible, but the government cannot allow an infinite net subsidy. The government solves the following problem, having decided on an upper bound for the net subsidy.
@@
()$(\ast)\begin{cases}\quad\max_{\{\tau_{t}\}_{t=0}^{\infty}}(1-\beta)\mathrm{E}_{0}\left[\sum_{t=0}^{\infty}\beta^{t}u(c_{t})\right]\\ \text{subject to}\\ \quad(1-\beta)\mathrm{E}_{0}\left[\sum_{t=0}^{\infty}\beta^{t}(\tau_{t})\right]\leq\text{Const},\end{cases}$
@@
where $\mathrm{E}_{0}$ is the conditional expectation given $e_{0}$, $u:\mathbb{R}_{+}\to\mathbb{R}$ is the individual’s utility function, $c_{t}=e_{t}+\tau_{t}$ is the consumption in period $t$ starting from the individual’s birth period. The utility function $u$ is assumed to be $u^{\prime}>0$, $u^{\prime\prime}<0$, to satisfy the Inada conditions ($\lim_{c\to 0}u^{\prime}(c)=+\infty$, $\lim_{c\to\infty}u^{\prime}(c)=0$) and to be bounded from above.
@@
Without specifying the upper bound of the net subsidy, $\{\tau_{t}\}_{t=0}^{\infty}$ is determined by $\lambda_{0}$, and the same $\lambda_{0}$ determines the upper bound of the net subsidy as well. Thus, the problem ()($\ast$)( ∗ ) can be rewritten as the following problem, where the objective function is the Lagrangian of problem ()($\ast$)( ∗ ) with an undetermined multiplier $01/\lambda_{0}>$ 0.
$\max_{\{\tau_{t}\}_{t=0}^{\infty}}$ $(1-\beta)\mathrm{E}_{0}\left[\sum_{t=0}^{\infty}\beta^{t}[\lambda_{0}u(c_{t})-\tau_{t}]\right].$ (1)
@@
and we obtain $c_{t}=(u^{\prime})^{-1}(\lambda_{0}^{-1})$ and $\tau_{t}=(u^{\prime})^{-1}(\lambda_{0}^{-1})-e_{t}$ for $t\in\mathbb{N}\cup\{0\}$. This is the Pareto optimal transfer mechanism that gives the individual constant consumption in each period.
Let $v_{1}:\mathbb{R}_{++}\times\mathbb{R}_{+}\to\mathbb{R}$ be a function expressing each individual’s lifetime utility provided by the transfer mechanism. Let $v_{2}:\mathbb{R}_{++}\times\mathbb{R}_{+}\to\mathbb{R}$ be a function expressing the expected balance of transfers. Then, for given $\lambda_{0}>0$ and $e_{0}\in E$, these values are written as follows.
$v_{1}(\lambda_{0},e_{0})$ $:=(1-\beta)\mathrm{E}_{0}[\sum_{t=0}^{\infty}\beta^{t}u(c_{t})]=u((u^{\prime})^{-1}(\lambda_{0}^{-1}))=:\overline{v}_{1}(\lambda_{0})$
$v_{2}(\lambda_{0},e_{0})$ $:=(1-\beta)\mathrm{E}_{0}[\sum_{t=0}^{\infty}\beta^{t}(-\tau_{t})].$
Note that the function $v_{1}$ could be seen as a function of the value $\lambda_{0}$ whose inverse was the Lagrangian multiplier. We write it as $\overline{v}_{1}$. The function $\overline{v}_{1}$ is bijective and increasing. On the other hand, we see that $v_{2}(\lambda_{0},e_{0})$ depends on income. So we write its expectation as $\overline{v}_{2}(\lambda_{0})$.
$\overline{v}_{2}(\lambda_{0})$ $:=\mathrm{E}[v_{2}(\lambda_{0},e_{t})].$
In Section 4 the government will adjust the value of $\lambda_{0}$ so that the expected balance of total transfers across generations is non-negative.
Let $\tau^{*}(\cdot\ |\lambda_{0})$ be a time-invariant transfer function, which is the solution to the problem (1). That is, $\tau^{*}:\mathbb{R}_{+}\times\mathbb{R}_{++}\to\mathbb{R}_{+}$ is defined such that, for a reported income $x\in\mathbb{R}_{+}$ and a utility control variable $\lambda>0$
$\tau^{*}(x|\lambda):=(u^{\prime})^{-1}(\lambda^{-1})-x.$
We will use these notation in the following section for the case where income is private information. There we observe that a utility control variable $\lambda_{t}$ can play the role of a state variable and the role of an incentive device. Let $\mathcal{T}$ be the set of all admissible control functions there, and let $\tau:\mathbb{R}_{+}\times\mathbb{R}_{++}\to\mathbb{R}_{+}$ be a typical element of $\mathcal{T}$. Then the lifetime utility of individuals who are guaranteed utility level $\lambda_{t}$ at an initial period $t$ is given by
$J(\lambda_{t},t,\tau)=(1-\beta)\mathrm{E}_{t}\left[\sum_{k=t}^{\infty}\beta^{k-t}u(e_{k}+\tau(e_{k}|\lambda_{t}))\right]\quad\text{for }\tau\in\mathcal{T},$ (2)
where $\mathrm{E}_{t}$ denotes the conditional expectation given $e_{t}$. We will see that $\tau^{*}\in\mathcal{T}$ could be an optimal control function for any initial period $t$ and for later specified $\{\lambda_{k}\}_{k=t}^{\infty}$ such that
$J(\lambda_{t},t,\tau^{*})=\sup_{\tau\in\mathcal{T}}$ $(1-\beta)\mathrm{E}_{t}\left[\sum_{k=t}^{\infty}\beta^{k-t}u(e_{k}+\tau(e_{k}|\lambda_{t}))\right].$ (3)
2.2 Pareto optimal transfer mechanism with asymmetric information
If income is private information, the government has to ask individuals to declare their income. In general, a mechanism $\Gamma$ that arranges transfers $\{\tau_{t}\}_{t=0}^{\infty}$ needs to satisfy sequential incentive compatibility defined below. For this definition, let $z$ denote the current observable state, which can be a finite vector of state variables, and may include the past history of observables or be in a recursive form. The distribution of the tomorrow’s state $z^{\prime}$ is determined by this $z$ and the current level of consumption $x$. We write $\pi(x)$ for the value of each individual’s current payoff and $v(z)$ for the present value of the mechanism $\Gamma$ at $z$.
@@ 2.2 Pareto optimal transfer mechanism with asymmetric information
[Sequential incentive compatibility, Marcet and Marimon, 1992] The mechanism $\Gamma$ is said to be sequentially incentive compatible if for every sate $z$
$v(z)=(1-\beta)\pi(x)+\beta\mathrm{E}_{(x,z)}[v(z^{\prime})]\geq(1-\beta)\pi(\tilde{x})+\beta\mathrm{E}_{(\tilde{x},z)}[v(z^{\prime})],$
where $E_{(x,z)}$ means that the conditional expectation given $(x,z)$.
@@
[Sequentially efficient mechanism, Marcet and Marimon, 1992] The mechanism $\Gamma$ is said to be sequentially efficient mechanism if it is sequentially incentive compatible and it is not Pareto dominated by any other sequnetially incentive compatible mechanism.
In addition to taking these things into account, since it is known to lead to an immisarization outcome, the government does not hope to manipulate the transfer function $\tau^{*}$. The government takes an alternative approach. It is to vary the utility control variable $\lambda$ in each period. This could be an incentive to tell the truth. For a given $\lambda$ in a given period, the government promises to ensure that the lifetime utility level $\overline{v}_{1}(\lambda)$, and the utility control variable for the next period is renewed taking into account the current utility control variable and the income reported in the current period. If an individual truthfully reports income $e$ in a period $t$, then the lifetime utility from that period should be
$\overline{v}_{1}(\lambda)=\mathrm{E}\big{[}(1-\beta)u(e+\tau^{*}(e|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(e|\lambda))\big{]},$ (4)
where $\lambda^{\prime}(\cdot|\lambda):E\to\mathbb{R}_{++}$ is the utility control variable for the next period. If $\lambda^{\prime}$ satisfies (4), then $\mathrm{E}[\overline{v}_{1}(\lambda^{\prime}(e|\lambda)]=\overline{v}_{1}(\lambda)$. Truth telling is then induced if, given the current utility control variable $\lambda$, the renewed utility control variable $\lambda^{\prime}$ satisfies the following incentive constraints.
@@
Such an utility control variable $\lambda^{\prime}$ that satisfies both (4) and (5) could be adapted from the $\lambda$ mechanism of Marcet and Marimon [1992], provided that there is a feasible $\lambda^{\prime}$. By a feasible $\lambda^{\prime}$ we mean that the updated utility control variable keeps the level of expected utility for the next period below the supremum of each individual’s utility function. That is, let $S$ be the supremum of each individual’s utility function, $S=\sup_{x\in\mathbb{R}_{+}}u(x)$. For given $\lambda>0$, if $\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,e^{M})-\overline{v}_{2}(\lambda))<S$, define $\lambda^{\prime}$ such that
@@
Since $v_{2}(\lambda,e)-\overline{v}_{2}(\lambda)=(1-\beta)(e-\mathrm{E}[e_{t}])$, if the reported income $e\in E$ is above the average income, $\mathrm{E}[e_{t}]$, this will be rewarded with higher utility control variables for the next period. Conversely, if the reported income $e\in E$ is below the average income, $\mathrm{E}[e_{t}]$, this will result in lower utility control variables for the next period. Note that $u(e+\tau^{*}(e|\lambda))=u((u^{\prime})^{-1}(\lambda^{-1}))=\overline{v}_{1}(\lambda)$. From this we see that $\lambda^{\prime}$ defined in (6) satisfies (4):
$\mathrm{E}[(1-\beta)u(e+\tau^{*}(e|\lambda))$ $+\beta\overline{v}_{1}(\lambda^{\prime}(e|\lambda))]$
$=\mathrm{E}[\overline{v}_{1}(\lambda)+\lambda^{-1}(v_{2}(\lambda,e)-\overline{v}_{2}(\lambda))]=\overline{v}_{1}(\lambda).$
Furthermore, remind that the government’s objective function in problem (1) with $\lambda_{0}=\lambda$:
$\lambda\big{[}(1-\beta)u(e+\tau^{*}(e|\lambda))+\beta\overline{v}_{1}(\lambda)\big{]}+v_{2}(\lambda,e)$
is optimal at true income report $e$. It follows that
$\lambda\big{[}(1-\beta)$ $u(e+\tau^{*}(e|\lambda))+\beta\overline{v}_{1}(\lambda)\big{]}+v_{2}(\lambda,e)-\overline{v}_{2}(\lambda)$
$\geq\lambda[(1-\beta)u(e+\tau^{*}(\tilde{e}|\lambda))+\beta\overline{v}_{1}(\lambda)]+v_{2}(\lambda,\tilde{e})-\overline{v}_{2}(\lambda)\quad\text{ for }e,\tilde{e}\in E.$
Then we have the following inequality and see that (5) holds.
$\lambda\big{[}(1-\beta)u(e+\tau^{*}(e|\lambda))$ $+\beta\overline{v}_{1}(\lambda)\big{]}+v_{2}(\lambda,e)-\overline{v}_{2}(\lambda)$
$=\lambda\big{[}(1-\beta)u(e+\tau^{*}(e|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(e|\lambda))\big{]}$
$\geq\lambda[(1-\beta)u(e+\tau^{*}(\tilde{e}|\lambda))+\beta\overline{v}_{1}(\lambda)]+v_{2}(\lambda,\tilde{e})-\overline{v}_{2}(\lambda)$
$=\lambda\big{[}(1-\beta)u(e+\tau^{*}(\tilde{e}|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(\tilde{e}|\lambda))\big{]}\quad\text{ for }e,\tilde{e}\in E.$
On the other hand, if there is an income $\hat{e}\in E$ for given $\lambda$ such that $\lambda^{\prime}$ defined in (6) must promise the expected utility for the next period above the supremum of the individual’s utility function, then the government must find the closest utility control variable for which there is a certainty equivalent to the promised level of lifetime utility. That is, for given $\lambda$, if there is $\hat{e}\in E$ such that $\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,\hat{e})-\overline{v}_{2}(\lambda))\geq S$, for reported income $e\in E$, define $\lambda^{\prime}$ such that
$\lambda^{\prime}(e|\lambda)=\overline{v}_{1}^{-1}\big{(}\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,\min\{e,\overline{e}_{\lambda}\})-\mathrm{E}[v_{2}(\lambda,\min\{e_{t},\overline{e}_{\lambda}\})])\big{)},$ (7)
where $\overline{e}_{\lambda}$ is defined such that
$\overline{e}_{\lambda}=\arg\max_{e^{\prime}\in E_{N}}\big{(}v_{2}(\lambda,e^{\prime})-\mathrm{E}[v_{2}(\lambda,\min\{e_{t},e^{\prime}\})]\big{)}$
@@
$\overline{v}_{1}(\lambda)+\frac{1}{\lambda\beta}\big{(}v_{2}(\lambda,e^{\prime})-\mathrm{E}[v_{2}(\lambda,\min\{e_{t},e^{\prime}\})]\big{)}<S,$
where $E_{N}\supset E$ is a set with $|E_{N}|=N>M$, $N\in\mathbb{N}$, and for $e\in E_{N}$ $e^{1}\leq e\leq e^{M}$. [^2] To avoid getting stuck on a pass where the same utility control variable $\lambda$ goes on forever, if you find $\overline{e}_{\lambda}=e^{1}$ for given $N$, choose $N^{\prime}\in\mathbb{N}$ with $N^{\prime}>N$, and use $E_{N^{\prime}}$ instead of $E_{N}$ so as to obtain $\overline{e}_{\lambda}\geq e^{2}$. This adjustment also changes the transfers in the current period $t$. However, we see the recursive relation in (7) satisfies the following corresponding incentive constraint.

### v15

@@ Abstract
We study dynamic mechanism design in a pure-exchange economy with privately observed idiosyncratic income. Classic hidden-income contracts attain constrained efficiency only at the cost of immiseration (Green 1987; Thomas–Worrall 1990). We propose a simple recursive mechanism—adapted from Marcet–Marimon (1992)—that shifts each income shock forward by one period, keeps promised utilities in a bounded set, and, under a transparent “moderate risk-aversion” condition, delivers sequential efficiency. In a stationary overlapping-generations setting, we further provide an explicit condition on the initial promise that ensures budget sustainability; early cohorts pre-fund intertemporal smoothing so that every cohort attains a higher expected lifetime utility than under autarky. Our analysis uses a single state (promised utility), closed-form transfers, and a Bellman verification.
Keywords: Dynamic mechanism design; recursive contracts; private information; immiseration; overlapping generations.
JEL: D82, D86.
@@ 1 Introduction
Private information can render idiosyncratic income effectively uninsurable. Since the seminal work of Green [1987] and Thomas and Worrall [1990], the constrained-efficient hidden-income contract trades insurance off against incentives and features a drift of average promised utilities toward a lower bound—immiseration—with cross-sectional dispersion rising over time. [^2] Earlier papers sometimes use immiserization[e.g., Phelan, 1998, Zhang, 2009]. Recent work with persistent private information, Bloedel et al. [2025], also sharpens the mechanism of immiseration and explains the balance of the cost of incentive provision and backloading of high-powered incentives. When agents are almost perfectly patient, the drift can vanish [Carrasco et al., 2019], but the knife-edge offers little comfort for moderate discounting.
A different route was proposed by Marcet and Marimon [1992]. Allowing agents to make productivity-enhancing investments, they designed a transfer rule that shifts risk across periods for the same individual. Capital accumulation enlarges the future consumption set and breaks the immiseration trap; the associated recursive saddle-point representation, later formalized by Marcet and Marimon [2019] makes the analysis more tractable. Yet the mechanism has not been explored in a pure-exchange economy without investment or storage, where the scope for avoiding downward drift is a priori unclear.
This paper adapts the Marcet–Marimon risk-shifting mechanism to an exchange economy in which agents cannot save. Promised utility is the sole state variable, so the model yields a closed-form characterization. If individual assets were allowed, one would require the set-valued approximation methods of Sleet and Yeltekin [2016]—a contrast confirmed by our closed-form example in Section 4.
The contribution of my paper is threefold. (i) A recursive, one-state dynamic mechanism that postpones shocks one period ahead and eliminates immiseration while preserving sequential incentive compatibility; (ii) a transparent sufficient condition for sequential efficiency under moderate risk aversion; (iii) in a stationary overlapping-generations society, an explicit inequality on the initial promise that suffices for budget sustainability, with a simple economic interpretation; early cohort pre-fund later transfers; the mechanism therefore delivers higher expected lifetime utility than autarky for all cohorts.
The organization of the paper is as follows. Section 2 sets up the model and transfer rule. Section 3 proves efficiency and boundedness, and analyzes inter-generational sustainability. Section 4 presents numerical illustrations. Section 5 verifies optimality via a Bellman equation, and Section 6 concludes.
@@ 2 Model
2.1 Environment
We consider a risk-neutral planner allocating transfers across a continuum of risk-averse individuals indexed by birth cohorts. Individuals survive to the next period with probability $\alpha\in(0,1)$ and discount at rate $r\in(0,1)$; their effective discount factor is $\beta=\alpha r$. In each period, a newborn cohort enters so that the cross-section is stationary. [^3] A similar stationary population model is used in Fujiwara-Greve and Okuno-Fujiwara [2009] for their Prisoner’s Dilemma game. Each individual draws an i.i.d. income shock $e_{t}\in E:=\{e^{1},e^{2},\dots,e^{M}\}\subset\mathbb{R}_{+}$ with $e^{i}<e^{j}$ for $i<j\in\{1,\dots,M\}$, $M\geq 2$. The planner commits to a recursive transfer mechanism $\mathcal{M}$ mapping the current promised $\lambda_{t}>0$ into a transfer $\tau_{t}$ and next period’s promise $\lambda_{t+1}$. The single state variable is the promised utility $\overline{v}_{1}(\lambda)$, which is strictly increasing in $\lambda$.
2.2 Symmetric-information benchmark
@@
First, we derive an optimal transfer function which, under symmetric income information, ensures a level of utility corresponding to a promise $\lambda_{0}>0$ . Following Marcet and Marimon [1992], we will use this promise $\lambda_{0}$ as a state variable under asymmetric income information. For this purpose, in this subsection only, we consider the case where individuals cannot hide their income from the planner in each period.
@@
At period 0, the planner is considering an efficient transfer mechanism that increases individuals’ lifetime utility as much as possible, but it cannot afford an infinite net subsidy. The planner solves the following problem, which is normalized by multiplying $(1-\beta)$, having decided on an upper bound for the net subsidy.
@@
$(\ast)\quad\max_{\{\tau_{t}\}_{t=0}^{\infty}}(1-\beta)\mathbb{E}_{0}\left[\sum_{t=0}^{\infty}\beta^{t}u(c_{t})\right]\quad\text{s.t. }\quad(1-\beta)\mathbb{E}_{0}\left[\sum_{t=0}^{\infty}\beta^{t}(\tau_{t})\right]\leq\overline{C},$
@@
where $\overline{C}>0$ is an exogenous upper bound on the present value of net subsidies, $\mathbb{E}_{0}$ is the conditional expectation given the information in period 0, $u:\mathbb{R}_{+}\to\mathbb{R}$ is the individual’s utility function, $c_{t}=e_{t}+\tau_{t}$ is the consumption in period $t$ starting from the individual’s birth period. The utility function $u$ is assumed to be $u^{\prime}>0$, $u^{\prime\prime}<0$, to satisfy the Inada conditions ($\lim_{c\to 0}u^{\prime}(c)=+\infty$, $\lim_{c\to\infty}u^{\prime}(c)=0$) and to be bounded from above.
@@
Without specifying $\overline{C}$, $\{\tau_{t}\}_{t=0}^{\infty}$ is determined by $\lambda_{0}$, and the same $\lambda_{0}$ determines $\overline{C}$ as well. Let $\mu>0$ be the multiplier on the subsidy constraint and put $\lambda_{0}=1/\mu$. The problem $(\ast)$ can be rewritten as the following problem.
$\max_{\{\tau_{t}\}_{t=0}^{\infty}}$ $(1-\beta)\mathbb{E}_{0}\left[\sum_{t=0}^{\infty}\beta^{t}[\lambda_{0}u(c_{t})-\tau_{t}]\right].$ (1)
@@
and the optimal policy delivers a constant consumption $c^{\ast}(\lambda_{0})=(u^{\prime})^{-1}(\lambda_{0}^{-1})$ and transfer $\tau^{\ast}(e|\lambda_{0})=(u^{\prime})^{-1}(\lambda_{0}^{-1})-e$, where $\tau^{*}(\cdot\ |\lambda_{0})$ is a time-invariant transfer function. We record two objectives for future use: Let $v_{1}:\mathbb{R}_{++}\times\mathbb{R}_{+}\to\mathbb{R}$ be a function expressing each individual’s lifetime utility provided by the transfer mechanism. Let $v_{2}:\mathbb{R}_{++}\times\mathbb{R}_{+}\to\mathbb{R}$ be a function expressing the expected balance of transfers. Then, for given $\lambda_{0}>0$ and $e_{0}\in E$, these values are written as follows.
$v_{1}(\lambda_{0},e_{0})$ $:=(1-\beta)\mathbb{E}_{0}[\sum_{t=0}^{\infty}\beta^{t}u(c_{t})]=u((u^{\prime})^{-1}(\lambda_{0}^{-1}))=:\overline{v}_{1}(\lambda_{0})\,\text{(promised utility)}$
$v_{2}(\lambda_{0},e_{0})$ $:=(1-\beta)\mathbb{E}_{0}[\sum_{t=0}^{\infty}\beta^{t}(-\tau_{t})]\,\text{(expected budget impact)}.$
The map $\lambda\mapsto\overline{v}_{1}(\lambda)$ is strictly increasing and will serve as the state variable once income becomes private. We also write
$\overline{v}_{2}(\lambda_{t}):=\mathbb{E}[(1-\beta)\mathbb{E}_{t}[\sum_{s=t}^{\infty}\beta^{s-t}(-\tau^{\ast}(e_{s}|\lambda_{s}))]].$
Remark 1.
(Lifetime value functional) Let $\mathcal{T}$ be the set of all admissible control functions. Given an initial promise $\lambda_{t}$ and any admissible transfer rule $\tau\in\mathcal{T}$, each individual’s discounted lifetime utility is
$J(\lambda_{t},t,\tau)=(1-\beta)\mathbb{E}_{t}\left[\sum_{k=t}^{\infty}\beta^{k-t}u(e_{k}+\tau(e_{k}|\lambda_{t}))\right]\quad\text{for }\tau\in\mathcal{T}.$ (2)
This function will be used in Section 5 to verify that our mechanism solves the associated Bellman equation.
2.3 Private information and mechanism design
If income is private information, the planner has to ask individuals to declare their income. In general, a mechanism $\Gamma$ that arranges transfers $\{\tau_{t}\}_{t=0}^{\infty}$ needs to satisfy sequential incentive compatibility defined below.
@@ 2.3 Private information and mechanism design
[Sequential incentive compatibility, Marcet and Marimon, 1992] Given the current promise $\lambda$, truthful reporting maximizes the continuation value. Equivalently,
$(1-\beta)u(e+\tau(e|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(e|\lambda))\geq(1-\beta)u(e+\tau(\tilde{e}|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(\tilde{e}|\lambda))\quad\forall e,\tilde{e}\in E.$
@@
[Sequentially efficient mechanism, Marcet and Marimon, 1992] The mechanism $\Gamma$ is said to be a sequentially efficient mechanism if it is sequentially incentive compatible and it is not Pareto dominated by any other sequentially incentive compatible mechanism.
In addition to taking these things into account, since it is known to lead to an immiseration outcome, the planner does not hope to manipulate the transfer function $\tau^{*}$. The planner takes an alternative approach. It is to vary the promise $\lambda$ in each period. This could be an incentive to tell the truth. For a given $\lambda$ in a given period, the planner promises to ensure that the lifetime utility level $\overline{v}_{1}(\lambda)$, and the promise for the next period is renewed taking into account the current promise and the income reported in the current period. If an individual truthfully reports income $e$ in a period $t$, then the lifetime utility from that period should be
$\overline{v}_{1}(\lambda)=\mathbb{E}\big{[}(1-\beta)u(e+\tau^{*}(e|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(e|\lambda))\big{]},$ (3)
where $\lambda^{\prime}(\cdot|\lambda):E\to\mathbb{R}_{++}$ is the promise for the next period. If $\lambda^{\prime}$ satisfies (3), then $\mathbb{E}[\overline{v}_{1}(\lambda^{\prime}(e|\lambda)]=\overline{v}_{1}(\lambda)$. Truth-telling is then induced if, given the current promise $\lambda$, the renewed promise $\lambda^{\prime}$ satisfies the following incentive constraints.
@@
Such a promise $\lambda^{\prime}$ that satisfies both (3) and (4) could be adapted from the $\lambda$ mechanism of Marcet and Marimon [1992], provided that there is a feasible $\lambda^{\prime}$. By a feasible $\lambda^{\prime}$ we mean that the updated promise keeps the level of expected utility for the next period below the supremum of each individual’s utility function. That is, let $S$ be the supremum of each individual’s utility function, $S=\sup_{x\in\mathbb{R}_{+}}u(x)$. For given $\lambda>0$, if $\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,e^{M})-\overline{v}_{2}(\lambda))<S$, define $\lambda^{\prime}$ such that
@@
Under the updating rule of $\lambda^{\prime}$ in (5), misreporting never raises the weighted sum $\lambda u(e+\tau^{*}(e|\lambda))+v_{2}(\lambda,e)$, hence truth-telling is optimal.
On the other hand, if there is an income $\hat{e}\in E$ for given $\lambda$ such that $\lambda^{\prime}$ defined in (5) must promise the expected utility for the next period above the supremum of the individual’s utility function, then the planner must find the closest promise for which there is a certainty equivalent to the promised level of lifetime utility. That is, for given $\lambda$, if there is $\hat{e}\in E$ such that $\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,\hat{e})-\overline{v}_{2}(\lambda))\geq S$, for reported income $e\in E$, define $\lambda^{\prime}$ such that
$\lambda^{\prime}(e|\lambda)=\overline{v}_{1}^{-1}\bigg{(}\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,\min\{e,\overline{e}_{\lambda}\})-\mathbb{E}_{\tau^{\ast}}\left[v_{2}(\lambda,\min\{e,\overline{e}_{\lambda}\})\right]\bigg{)},$ (6)
where $\mathbb{E}_{\tau^{\ast}}$ denotes
$\mathbb{E}_{\tau^{\ast}}\left[v_{2}(\lambda,\min\{e,\overline{e}_{\lambda}\})\right]:=(1-\beta)\mathbb{E}\left[-\tau^{\ast}(\min\{e_{t},\overline{e}_{\lambda}\}|\lambda)+\sum_{s=t+1}^{\infty}\beta^{s-t}(-\tau^{\ast}(e_{s}|\lambda_{s}))\right]$
$\overline{e}_{\lambda}$ is defined such that
$\overline{e}_{\lambda}=\arg\max_{e^{\prime}\in E}\big{(}v_{2}(\lambda,e^{\prime})-\mathbb{E}_{\tau^{\ast}}[v_{2}(\lambda,\min\{e_{t},e^{\prime}\})]\big{)}$
@@
$\overline{v}_{1}(\lambda)+\frac{1}{\lambda\beta}\big{(}v_{2}(\lambda,e^{\prime})-\mathbb{E}_{\tau^{\ast}}[v_{2}(\lambda,\min\{e_{t},e^{\prime}\})]\big{)}<S.$
This adjustment also changes the transfers in the current period $t$. However, we see the recursive relation in (6) satisfies the following corresponding incentive constraint.